A random sample of eight observations from three normally distributed populations produced the

following data:

Treatments

A B C

84 78 87

79 74 80

87 81 91

85 86 77

94 86 78

89 89 79

89 69 77

83 79 78

Sample mean 86:25 80:25 80:88

Sample variance 20:79 45:07 27:27

Which one of the following statements is incorrect?

1. The grand mean is 82:46.

2. The sum squares for treatments .SST / is 173:9568.

3. The sum squares of error .SSE/ is 651:91:

4. The mean squares of error .M SE/ is 3:0433:

5. The F test statistic is 2:802:


1
Expert's answer
2021-08-20T11:47:56-0400

the grand mean

"X_{GM}=\frac{86.25+ 80.25 +80.88}{3}\\
=\frac{247.38}{3}\\
=82.46"

It is true


The sum squares

"SST= \sum n(x-\bar{x})\\
=8(82.4-82.46)^2+8(82.25-82.46)^2+8(80.88-82.46)^2\\
=8\cdot \:0.06^2+8\left(82.25-82.46\right)^2+8\left(80.88-82.46\right)^2\\
=8\cdot \:0.06^2+8\cdot \:0.21^2+8\cdot \:1.58^2\\
=20.3528"

It is incorrect


The sum squares of error

"SSE= \sum (n-1)s^2\\
=(8-1)*20.79^2+(8-1)*45.07^2+(8-1)*27.27^2\\
=7\cdot \:20.79^2+7\cdot \:45.07^2+7\cdot \:27.27^2\\
=3025.5687+14219.1343+5205.5703\\
=22450.2733"

It is incorrect


The mean squares of error

"MSE = \frac{SSE}{3}\\
= \frac{22450.2733}{3}\\
=7483.42443"

It is incorrect


The F test statistic

"F= \frac{SST/3-1}{SSE/30-3}\\
F= \frac{20.3528/3-1}{22450.2733/30-3}\\
=\frac{\frac{20.3528}{2}}{\frac{22450.2733}{30-3}}\\
=\frac{549.5256}{44900.5466}=0.01223"

It is incorrect